OpenStudy (anonymous):

The length of a car's skid mark in feet is given by L=0.046s2−0.199s+0.264,where sis the car's speed in miles per hour. If the length of a skid mark is 100feet, find the speed in miles per hour the car was traveling.

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@Jared25

OpenStudy (anonymous):

@1DEA

OpenStudy (anonymous):

\[L=0.046s ^{2}-0.199s+2.264\] correct?

OpenStudy (aum):

Solve for s: \[100 = 0.046s^2−0.199s+0.264\]

OpenStudy (anonymous):

yes, that is correct, but I am not sure how to solve it

OpenStudy (anonymous):

@aum @jared25

OpenStudy (aum):

\[ 100 = 0.046s^2−0.199s+0.264 \\ \text{Subtract 100:}\\ 0 = 0.046s^2−0.199s-99.736 \\ 0.046s^2−0.199s-99.736 = 0 \\ \text{Divide by 0.046:}\\ s^2 - 4.3261s - 2168.17 = 0 \\ s = \frac{4.3261 \pm \sqrt{(-4.3261)^2 -4*1*(-2168.17)}}{2*1} = \frac{4.3261 \pm 93.2277}{2} = 48.8, - 44.5 \\ s = 48.8 \text{ miles / hour}\]